# Constraints on inflationary decoherence from attractor model

### v1

@@ Abstract
The quantum-to-classical transition of inflationary perturbations remains a fundamental challenge in cosmology, and quantum decoherence is one of the proposed solutions to this issue. By treating primordial quantum perturbations as an open quantum system interacting with environmental degrees of freedom during inflation, decoherence can be described by the Lindblad master equation. Lindblad equation modifies the dynamical evolution of the primordial quantum perturbations and affects the power spectrum of the curvature perturbations, resulting in changes in the primordial power spectrum. In this paper, we examine the decoherence process of a polynomial attractor model that includes an ultra-slow-roll plateau. We compare our findings with previous results from slow-roll analyses. We numerically compute the corrections to the power spectrum due to decoherence. The results show significant modifications only on large scales, with a peak appearing at the minimum of the power spectrum. Using CMB constraints on the spectral index and tensor-to-scalar ratio, and requiring complete decoherence for relevant scales by the end of inflation, we constrain the interaction parameter: $10^{-28}\text{Mpc}^{-1}<k_{\gamma}<0.042\text{Mpc}^{-1}$.
@@ I Introduction
The inflationary theory is one of the most successful theories in cosmology today. It assumes that the universe experienced a rapid expansion in its very early stages, which naturally addresses the horizon and flatness problems of the universe [1, 2, 3, 4, 5]. The primordial quantum fluctuations during the inflationary epoch can provide the primordial density perturbations that lead to the formation of large-scale structures (LSS) in the universe and anisotropies in the Cosmic Microwave Background (CMB) radiation [6, 7, 8, 9, 10, 11], which is supported by observations [12].
@@
However, inflationary theory faces several unresolved challenges, with one of the most important being the quantum-to-classical transition problem [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. The observational quantities in modern cosmology, including the temperature anisotropies of the CMB and the statistical properties of the LSS, demonstrate classical behavior. The question is how the primordial quantum fluctuations that seeded these structures made the transition to classical ones. A central issue arises from modeling primordial quantum perturbations as isolated systems. In reality, these perturbations interact with environmental degrees of freedom, leading to decoherence [38, 39, 40] that effectively converts quantum fluctuations into classical stochastic density inhomogeneities.
The dynamics of an open quantum system are described by the Lindblad master equation [41], which describes the time evolution of a quantum system interacting with an external environment. Under the environmental interactions, decoherence suppresses the off-diagonal elements of the system’s density matrix, driving it to a diagonal state [42, 43, 44]. This equation leads to a correction to the outcomes of the primordial perturbations, thereby affecting their power spectrum. In this paper, we consider the effect of decoherence and numerically compute the modified power spectrum of a inflationary model with an ultra-slow-roll stage. Moreover, we constrain the interaction parameter from the CMB observations.
The process of decoherence takes time and is considered complete when the off-diagonal elements of the quantum system’s density matrix become negligible. The decoherence time depends on the strength of the interaction between the system and its environment: stronger coupling leads to shorter timescales. On cosmological scales, the primordial quantum perturbations undergo complete decoherence by the end of inflation, thereby transitioning to a classical state. In this paper, we require that decoherence occurs sufficiently at the scale of interest, which imposes a lower bound constraint on the interaction parameter.
@@
This paper is organized as follows. In Sec. II, we present the Lindblad equation that the quantum perturbations satisfy during inflation, along with its exact solution under the assumption of linear interactions. In Sec. III, we provide the parameters of our model and present the results of numerical calculations regarding the impact of decoherence, as well as constraints on the interaction strength. We summarize this paper in Sec. IV.
@@ II Lindblad equation and the modified power spectrum
where $\hat{v}_{k}$ is the Mukhanov-sasaki variable in Fourier space and $\hat{p}_{k}$ is its conjugate momentum, which obeys $\hat{p}_{k}=\hat{v}_{k}^{\prime}$. A prime denotes a derivative with respect to conformal time $\eta$. The frequency of mode $\bm{k}$ is given by
@@
where $\epsilon_{1}$ is the first slow-roll parameter, $\epsilon_{1}=-\dot{H}/H^{2}$, with $H$ representing the Hubble parameter and a dot indicating differentiation with respect to time. The Hamiltonian (1) describes a isolated system. However, the primordial perturbations could interact with other degrees of freedom in the early universe, making it more realistic to consider the perturbations as an open quantum system [32, 35].
For an open quantum system that interacts with the environment, we can describe the total system in the Hilbert space $\mathcal{E}=\mathcal{E}_{S}\otimes\mathcal{E}_{E}$, where the total Hamiltonian is
@@
where $H_{S}$ is the Hamiltonian of the system, which acts in $\mathcal{E}_{S}$, and $H_{E}$ is the Hamiltonian of the environment, which acts in $\mathcal{E}_{E}$, $H_{\mathrm{int}}$ is the interaction Hamiltonian and $g$ is the coupling constant. Suppose that the system and the environment only couple through local interactions, then the interaction Hamiltonian can be written as
@@
where $\hat{A}$ acts as a local operator within the system’s subsystem, while $\hat{B}$ represents a local operator associated with the environmental component. The symbol $\bm{x}$ represents the conformal spatial coordinate. According to the reasons mentioned in [32, 35], we can consider $\hat{A}$ to have the form
@@
where $n$ is a constant. By tracing out the environmental degrees of freedom from the total density matrix $\hat{\rho}_{\text{tot}}$, we can derive the reduced density matrix for the system $\hat{\rho}_{S}$
@@
Using the Born and Markov approximations, we can derive the equation of motion of $\hat{\rho}_{S}$, i.e. the Lindblad equation [41, 42, 32]
@@
where $\gamma=2g^{2}\eta_{c}$ with $\eta_{c}$ being the conformal auto-correlation time of $\hat{B}$, and $C_{B}$ is the same-time correlation function of $\hat{B}$, given by $C_{B}(\bm{x},\bm{y})=\text{Tr}_{E}[\rho_{E}\hat{B}(\eta,\bm{x})\hat{B}(\eta,\bm{y})]$. We assume that the time-dependent parameter $\gamma$ depends on the scale factor as a power law [32]
@@
where * denotes a reference time when the pivot scale $k_{*}=0.05\mathrm{Mpc}^{-1}$ crosses the Hubble radius, and $p$ is a free parameter. Assuming the environment to be isotropic and statistically homogeneous, $C_{B}(\bm{x},\bm{y})$ can be defined as a top hat function [32]
@@
Expanding Eq. (9) in Fourier space, we can obtain
@@
Considering the case where the environment consists of a heavy scalar field $\varphi$ with mass $M\gg H$, and the interaction Hamiltonian is
@@
where the parameter $\mu$ is defined as a constant mass scale and $\hat{\phi}=\hat{v}/a$. In this case, the parameters can be represented as [32]
@@
$\sigma(m)=\begin{cases}1,\quad m\,\,\mathrm{is\,\,even}\\ 0,\quad m\,\,\mathrm{is\,\,odd}\end{cases}.$ (16)
In order to solve the Lindblad equation (7), we decompose $\hat{v}_{\bm{k}}$ and $\hat{p}_{\bm{k}}$ into their real and imaginary parts
@@
For a real operator $\hat{v}(\eta,\bm{x})$, we have $\hat{v}_{-\bm{k}}=\hat{v}^{\dagger}_{\bm{k}}$, which implies $\hat{v}^{\mathrm{R}}_{-\bm{k}}=\hat{v}^{\mathrm{R}}_{\bm{k}}$ and $\hat{v}^{\mathrm{I}}_{-\bm{k}}=-\hat{v}^{\mathrm{I}}_{\bm{k}}$. These relations show that not all $\hat{v}_{\bm{k}}$ operators are independent degrees of freedom, we can calculate the variables $v_{\bm{k}}^{\mathrm{R}}$ and $v_{\bm{k}}^{\mathrm{I}}$ in $\bm{k}\in\mathbb{R}^{3+}$. Under the condition of linear interaction, the free Hamiltonian (1) can be written as
$\hat{H}_{\bm{k}}^{s}=\frac{1}{2}\int_{\mathbb{R}^{3+}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})(\hat{v}_{\bm{k}}^{s})^{2}+(\hat{p}_{\bm{k}}^{s})^{2}\right],$ (18)
@@
then we can obtain the Lindblad equation in the Fourier space for linear interaction [32]
@@
Note that Eq. (20) is valid only for linear interaction. We can define a combined interaction parameter as follows
@@
Then we define the eigenvectors $|v_{\rm k}^{s}\rangle$ of the operator $\hat{v}_{\rm k}^{s}$, which satisfy equation $\hat{v}_{\rm k}^{s}|v_{\mathbf{k}}^{s}\rangle=v_{\rm k}^{s}|v_{\rm k}^{s}\rangle$. By applying the bra $\langle v_{\bm{k}}^{s,(1)}|$ and the ket $|v_{\bm{k}}^{s,(2)}\rangle$ to the Lindblad equation Eq. (20), the Lindblad equation can be solved exactly, the solution is given by [32]
$\left\langle v_{\bm{k}}^{s,(1)}\right|\hat{\rho}_{\bm{k}}^{s}\left|v_{\bm{k}}^{s,(2)}\right\rangle=\frac{\left(2\pi\right)^{-1/2}}{\sqrt{\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}}}\mathrm{exp}\left\{-\frac{{v^{s,(2)}_{\bm{k}}}^{2}+{v^{s,(1)}_{\bm{k}}}^{2}+i{\left|v_{\bm{k}}\right|^{2}}^{\prime}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]}{4\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\right\}$
×$\qquad\qquad\times\exp\Biggl{\{}-\frac{\left[v^{s,(2)}_{\bm{k}}-v^{s,(1)}_{\bm{k}}\right]^{2}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left(\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\right)$
$\qquad\qquad-\frac{i\mathcal{K}_{\bm{k}}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]\Biggr{\}},$ (22)
where $\mathcal{I}_{\bm{k}}$, $\mathcal{J}_{\bm{k}}$ and $\mathcal{K}_{\bm{k}}$ are defined by
@@
where $v_{\bm{k}}(\eta)$ is the solution of the Mukhanov-Sasaki equation. Using Eq. (II) and the relation $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{\rho}_{v}\hat{O})$ with $\hat{O}=\hat{v}_{\bm{k}}^{2}$, we can obtain the two-point correlation function
@@
where we take $M_{\mathrm{Pl}}^{2}=1$. On the other hand, for linear interaction, the tensor power spectrum remains unchanged by the corrections in the Lindblad equation [35]. Therefore, we can utilize the spectral index and the tensor-to-scalar ratio to constrain $k_{\gamma}$ for linear interaction.
The interaction with the environment tends to suppress the off-diagonal elements of the density matrix when expressed in the basis of the eigenstates of the interaction operator [42, 32]. From the exact solution of the density matrix Eq. (II), we can study decoherence during inflation. The decoherence parameter $\delta_{\bm{k}}$ defined in [32] represents the reduction in the off-diagonal elements of the density matrix due to environmental interactions, and is defined as follows
@@
It is obvious that $\delta_{\bm{k}}$ is related to $k_{\gamma}$. When $\delta_{\bm{k}}\gg 1$, decoherence is expected to be complete. We determine that by the end of inflation, the decoherence of the scales of interest should be finished, which in turn constrains $k_{\gamma}$.
In the study referenced as [32], the researchers took into account the constant $\epsilon_{1}$ and the constant Hubble parameter $H$. In this paper, we will compute the results for a concrete model with an ultra-slow-roll stage. We compare the results of our model’s numerical computations with the analytical results reported in [32]. Then, we constraint $k_{\gamma}$ using the polynomial attractor model.

### v2

# Observational constraints on inflationary decoherence with polynomial attractor model

@@ Abstract
The quantum-to-classical transition of inflationary perturbations remains an unresolved fundamental problem, and quantum decoherence is one of the promising solutions. By considering quantum perturbations during inflation as an open quantum system interacting with its environment, quantum decoherence can be described by the Lindblad equation. This formalism modifies the evolution of primordial quantum perturbations and consequently alters the power spectrum of curvature perturbations, leading to observable consequences. In this paper, we examine the decoherence process of a polynomial attractor model featuring an ultra-slow-roll stage, extending previous analyses limited to slow-roll scenarios. We numerically compute the correction to the power spectrum due to quantum decoherence, and the results show significant modification only on large scales, with a peak generated by the decoherence correction at the minimum of the power spectrum. Using observational constraints on the scalar spectral index and the tensor-to-scalar ratio, and requiring complete decoherence for relevant scales by the end of inflation, we obtain the constraint on the interaction parameter as $10^{-17}\text{Mpc}^{-1}<k_{\gamma}<0.061\text{Mpc}^{-1}$.
@@ I Introduction
The inflationary scenario is one of the most successful progresses in cosmology. It assumes that the universe experienced a rapid expansion in its very early stages, which naturally addresses the monopole, horizon and flatness problems [1, 2, 3, 4, 5]. Quantum fluctuations during inflation generate primordial density perturbations that seed the formation of large-scale structures (LSS) observed today and imprint temperature anisotropies in the cosmic microwave background (CMB) radiation [6, 7, 8, 9, 10, 11].
@@
However, the inflationary scenario faces several unresolved challenges, with a significant one being the quantum-to-classical transition problem [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. The observables such as the temperature anisotropies of the CMB and the statistical properties of the LSS, are inherently classical in nature. This raises the critical question of how the primordial quantum fluctuations that seeded these structures became classical. The core issue stems from conventional treatments that model primordial perturbations as isolated quantum systems. In reality, these perturbations interact with environmental degrees of freedom, leading to decoherence [38, 39, 40] that effectively converts quantum fluctuations into classical stochastic density inhomogeneities.
The dynamics of an open quantum system are described by the Lindblad master equation, which governs the time evolution of a quantum system interacting with its environment. Such environmental couplings induce decoherence, causing the off-diagonal elements of the system’s density matrix to diminish and driving the density matrix towards a diagonal form [41, 42, 43]. This decoherence modifies primordial perturbations, leading to corrections in their power spectrum. The decoherence timescale—over which the off-diagonal elements become negligible—depends on the interaction strength between the system and its environment: stronger coupling results in faster decoherence. In this paper, we impose the condition that primordial quantum perturbations at relevant scales undergo complete decoherence by the end of inflation, ensuring that these perturbations become classical thereafter. We analyze the impact of decoherence through numerical computation of the modified power spectrum for a polynomial attractor model featuring an ultra-slow-roll stage. Furthermore, we derive constraints on the interaction parameter by employing the observational data P-ACT-LB-BK18 from the combination of Planck data, Atacama Cosmology Telescope (ACT) data, CMB lensing, baryon acoustic oscillation (BAO) distance measurements from Dark Energy Spectroscopic Instrument (DESI) and B-mode measurements from the BICEP/Keck telescopes (BK18) [44, 45].
@@
This paper is organized as follows. In Sec. II, we present the Lindblad equation along with its exact solution under the assumption of linear interaction. In Sec III, we present the results of numerical calculations regarding the impact of decoherence, as well as constraints on the interaction strength. The paper is concluded in Sec. IV.
@@ II Lindblad equation and the modified power spectrum
where $\hat{v}_{k}$ is the Mukhanov-sasaki variable in Fourier space, $\hat{p}_{k}=\hat{v}_{k}^{\prime}$ is its conjugate momentum, and a prime denotes a derivative with respect to conformal time $\eta$. The frequency of mode $\bm{k}$ is given by
@@
[…]
For an open quantum system that interacts with its environment, the total Hamiltonian is
@@
where $\hat{H}_{S}$ is the Hamiltonian of the quantum system which acts in the Hilbert space $\mathcal{E}_{S}$, $\hat{H}_{E}$ is the Hamiltonian of the environment which acts in the Hilbert space $\mathcal{E}_{E}$, $\hat{H}_{\mathrm{int}}$ is the interaction Hamiltonian and $g$ is the coupling constant. Suppose that the quantum system and its environment couple through local interactions only, then the interaction Hamiltonian can be written as
@@
where $\hat{A}$ acts as a local operator within the system’s subsystem, while $\hat{B}$ represents a local operator associated with the environmental component. Following Refs. [31, 34], we consider $\hat{A}$ to have the form
@@
where $n$ is a constant. By tracing out the environmental degrees of freedom from the total density matrix $\hat{\rho}_{\text{tot}}$, the reduced density matrix for the quantum system $\hat{\rho}_{S}$ becomes
@@
Using the Born and Markov approximations, the Lindblad equation for the reduced density matrix $\hat{\rho}_{S}$ is [46, 41, 31]
@@
where $\gamma=2g^{2}\eta_{c}$ with $\eta_{c}$ being the conformal auto-correlation time of $\hat{B}$, and $C_{B}(\bm{x},\bm{y})=\text{Tr}_{E}[\rho_{E}\hat{B}(\eta,\bm{x})\hat{B}(\eta,\bm{y})]$ is the same-time correlation function of $\hat{B}$. The time-dependent parameter $\gamma$ is assumed to vary with the scale factor following a power-law form [31],
@@
where * denotes a reference time when the pivot scale $k_{*}=0.05\text{ Mpc}^{-1}$ crosses the horizon, and $p$ is a free parameter. Assuming the environment to be isotropic and statistically homogeneous, $C_{B}(\bm{x},\bm{y})$ can be written as a top hat function [31]
@@
Expanding Eq. (9) in Fourier space, we obtain
@@
Considering the case where the environment consists of a heavy scalar field $\varphi$ with mass $M\gg H$, the interaction Hamiltonian is [31]
@@
where the parameter $\mu$ is a constant mass scale and $\hat{\phi}=\hat{v}/a$. In this case, we have [31]
@@
$\sigma(m)=\begin{cases}1,\quad\text{if }m\,\,\text{is even,}\\ 0,\quad\text{if }m\,\,\text{is odd.}\end{cases}$ (16)
For the convenience of solving the Lindblad equation (7), we decompose the operators $\hat{v}_{\bm{k}}$ and $\hat{p}_{\bm{k}}$ into their real and imaginary parts,
@@
For a real operator $\hat{v}(\eta,\bm{x})$, we have $\hat{v}_{-\bm{k}}=\hat{v}^{\dagger}_{\bm{k}}$, which implies $\hat{v}^{\mathrm{R}}_{-\bm{k}}=\hat{v}^{\mathrm{R}}_{\bm{k}}$ and $\hat{v}^{\mathrm{I}}_{-\bm{k}}=-\hat{v}^{\mathrm{I}}_{\bm{k}}$. These relations show that not all $\hat{v}_{\bm{k}}$ are independent degrees of freedom, we need to calculate the variables $v_{\bm{k}}^{\mathrm{R}}$ and $v_{\bm{k}}^{\mathrm{I}}$ only for $\bm{k}\in\mathbb{R}^{3+}$. Under the condition of linear interaction, the free Hamiltonian (1) can be written as
$\hat{H}_{\bm{k}}^{s}=\frac{1}{2}\int_{\mathbb{R}^{3}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})(\hat{v}_{\bm{k}}^{s})^{2}+(\hat{p}_{\bm{k}}^{s})^{2}\right],$ (18)
@@
The Lindblad equation (7) in the Fourier space for linear interaction becomes [31]
@@
Note that Eq. (20) is valid only for linear interaction. For the convenience of discussion, we introduce a combined interaction parameter,
@@
To solve the Lindblad equation, we define the eigenvectors $|v_{\rm k}^{s}\rangle$ of the operator $\hat{v}_{\rm k}^{s}$ satisfying equation $\hat{v}_{\rm k}^{s}|v_{\mathbf{k}}^{s}\rangle=v_{\rm k}^{s}|v_{\rm k}^{s}\rangle$. Projecting the Lindblad equation Eq. (20) with the bra $\langle v_{\bm{k}}^{s_{1}}|$ and the ket $|v_{\bm{k}}^{s_{2}}\rangle$, we get the the solution [31]
$\left\langle v_{\bm{k}}^{s_{1}}\right|\hat{\rho}_{\bm{k}}^{s}\left|v_{\bm{k}}^{s_{2}}\right\rangle=\frac{\left(2\pi\right)^{-1/2}}{\sqrt{\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}}}\mathrm{exp}\left\{-\frac{\left({v^{s_{2}}_{\bm{k}}}\right)^{2}+\left({v^{s_{1}}_{\bm{k}}}\right)^{2}+i{\left|v_{\bm{k}}\right|^{2}}^{\prime}\left[\left({v^{s_{2}}_{\bm{k}}}\right)^{2}-\left({v^{s_{1}}_{\bm{k}}}\right)^{2}\right]}{4\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\right\}$
×$\qquad\qquad\times\exp\Biggl{\{}-\frac{\left[v^{s_{2}}_{\bm{k}}-v^{s_{1}}_{\bm{k}}\right]^{2}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left(\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\right)$
$\qquad\qquad-\frac{i\mathcal{K}_{\bm{k}}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left[\left({v^{s_{2}}_{\bm{k}}}\right)^{2}-\left({v^{s_{1}}_{\bm{k}}}\right)^{2}\right]\Biggr{\}},$ (22)
where $\mathcal{I}_{\bm{k}}$, $\mathcal{J}_{\bm{k}}$ and $\mathcal{K}_{\bm{k}}$ are defined as
@@
and $v_{\bm{k}}(\eta)$ is the solution of the Mukhanov-Sasaki equation. Using Eq. (II) and the relation $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{\rho}_{v}\hat{O})$ with $\hat{O}=\hat{v}_{\bm{k}}^{2}$, we can obtain the two-point correlation function
@@
where we take $M_{\mathrm{Pl}}^{2}=1$. For linear interaction, the tensor power spectrum remains unchanged by the corrections introduced in the Lindblad equation [34]. Therefore, we can utilize the spectral index and the tensor-to-scalar ratio to constrain $k_{\gamma}$ for linear interaction.
The interaction with the environment tends to suppress the off-diagonal elements of the density matrix when expressed in the basis of the eigenstates of the interaction operator [41, 31]. From the exact solution to Eq. (II) satisfied by the density matrix, we can study decoherence during inflation. The parameter $\delta_{\bm{k}}$ defined in [31] represents the reduction in the off-diagonal elements of the density matrix due to environmental interactions, and is defined as follows
@@
It is obvious that $\delta_{\bm{k}}$ is related to $k_{\gamma}$. When $\delta_{\bm{k}}\gg 1$, decoherence is expected to be fully established. The decoherence process associated with cosmological scales of interest must have been completed by the end of the inflation, thereby establishing constraint on the interaction parameter $k_{\gamma}$. In [31], analytical expressions for the corrections to the power spectrum were derived within the slow-roll approximation. In this paper, we numerically calculate the correction to the power spectrum for a polynomial attractor model featuring an ultra-slow-roll stage, and compare the result with the analytical result reported in [31].